The vapor pressure of pure benzene (C6H6) at 25°C is 94.4 torrwhile the vapor pressure of pure
chloroform (CHCl3) at 25°C is 172 torr. An ideal solution isformed when 50 g of chloroform is added to 50 g
of benzene. Calculate the vapor pressure of chloroform over thischloroform-benzene solution at 25°C.

a) 86 torr
b) 68 torr
c) 148 torr
d) 125 torr
e) 172 torr
Answer #1
   
          
        
    
    
    
  Number of moles of benzene , n = (50 /78) = 0.6410
   Number of moles of chloroform, N = (50 / 119.5) = 0.4184
    
         molefraction of chloroform, X = (0.4180) / (0.4180 +0.6410)
                                                        = 0.3947
         Pressure of Chloroform over the solution =  0.3947 x (pressure of pure chloroform)
                                                                        = 0.3947 x 172 torr
                                                                          =67.9 torr
          ∴ The answer is a)